OpenStudy (anonymous):

(5-i)^6 expressed in trigonometric form and De Moivre's theorem

OpenStudy (amoodarya):

5-i=root(5^2+1^2) exp(I theta) tan theta = -1/5

OpenStudy (anonymous):

Would theta end up as -11.31?

OpenStudy (amoodarya):

yes th=-11.309932

OpenStudy (anonymous):

So would it end up as 6623.9+16280.0i?

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

Thanks

OpenStudy (amoodarya):

your welcome you did it by your self